Use the following information to answer questions 1 - 3: An investigation was made to see the effect of reaction temperature X on percent conversion of a chemical process Y. Observations of Y were recorded for 5 different temperatures. X contains repeated values. (The data can be found in exercise 11.35 on page 422 of the textbook.) A simple linear regression was run and the following output was produced ANOVA Source Df SS MS Regression 828.78 828.78 Error 10 966.14 Lack-of-Fit 3 408.31 136.10 7 557.83 S2-79.69 Pure Error Total 111794.92 1. What value should replace “A” in the ANOVA table? (What is the F-value for Regression?) Round your answer to 2 decimal places. What value should replace B in the ANOVA table? (What is the F-value for Lack-of-Fit?) Round your answer to 2 decimal places. In the test for Lack-of-Fit at a 0.05 the hypotheses are Ho: the linear model is adequate vs. Ha: higher order terms are needed What is the critical value for this test? 2. 3.
Answer #1

1) A = 828.78/79.69 = 10.40005 = 10.4

2) B = 136.10/79.69 = 1.707868 = 1.71

3) The critical value fot this test is F0.05;3,7 = 4.346831